OpenStudy (anonymous):

Will give medals . Which of the following is a polynomial with roots 3, -3 , and 3 ? (3 points) Answer x3 + 3x2 + 9x + 27 x3 - 3x2 + 9x - 27 x3 + 9x2 + 9x + 81 x3 - 9x2 + 9x - 81

OpenStudy (amistre64):

(3-3)(-3+3)(3-3) r r r (x-3)(x+3)(x-3)

OpenStudy (amistre64):

3^3 = 27; and --+ = +

OpenStudy (anonymous):

wait so what would the answer be ?

OpenStudy (amistre64):

i would assume that it would be the one that has a + and a 27 in close proximity ...

OpenStudy (amistre64):

my other assumption being that I thought you needed help in determining the solution; not in being tutored on the subject.

OpenStudy (anonymous):

\[ (x-3)^2(x+3) = 27-9 x-3 x^2+x^3 \]

OpenStudy (anonymous):

its a +9x

OpenStudy (sasogeek):

you mis-typed one of your answers... the answer is the second one with +9x actually being -9x :) so you'd have what ffm typed. x^3-3x^2-9x+27